# RAISE: Enhancing Scientific Reasoning in LLMs via Step-by-Step Retrieval

### v1

@@ 4.5 Analysis of RAISE
We also qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 4. While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations. For instance, in questions involving chemical reactions or physical formulas, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts. These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
@@ 5 Related Works
LLMs for Scientific Reasoning.    Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks. Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge. Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs (Rein et al., 2024; Zhong et al., 2025; Zhang et al., 2025b; Wang et al., 2024a; Jiang et al., 2024). Many works enhance scientific reasoning capabilities of LLMs through domain-specific training (Taylor et al., 2022; Prabhakar et al., 2025; Zhang et al., 2024b), step-by-step reasoning (Rueda et al., 2025; Rein et al., 2024), or retrieval of external knowledge or tools (Ma et al., 2024; Zhong et al., 2025; Wellawatte et al., 2025; Li et al., 2025). Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
Step-wise Reasoning.    A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs. An influential early approach, Chain-of-Thought prompting Wei et al. (2022), introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent. This inspired methods such as Plan-and-Solve Wang et al. (2023), which emphasizes high-level planning before answering, and Step-Back Prompting Zheng et al. (2024), which encourages abstraction by prompting the model to reflect before solving. Least-to-Most prompting Zhou et al. (2023) extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
@@
Retrieval-Augmented Generation.    Retrieval-Augmented Generation (RAG) Lewis et al. (2020b) was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation Lewis et al. (2020b); Karpukhin et al. (2020); Izacard and Grave (2021); Khattab and Zaharia (2020).
@@ 6 Conclusion
Limitations
In our study, we primarily conducted experiments using relatively smaller language models, mainly due to computational resource constraints. As a result, the performance observed may not fully reflect the potential of larger-scale models. Future work could explore scaling up the model size to assess its impact on reasoning capabilities and overall performance.
For the retrieval process, we used only DPR as the retriever. While DPR has been widely recognized for its effectiveness, we believe that more recent retrieval models Gao et al. (2023b); Basu et al. (2024); Shi et al. (2023), could lead to a substantial improvement in performance. Incorporating these advanced retrievers remains a promising direction for future research.
Furthermore, our work is particularly focused on scientific reasoning tasks. While this domain presents unique challenges and complexities, extending our approach to other domains such as legal, commonsense, or medical reasoning could be an interesting and valuable avenue for further investigation. Such cross-domain studies may help validate the generalizability and robustness of RAISE.
Additionally, RAISE relies on multiple calls to large language models during the logical query generation phase. Although this multi-step querying is essential for generating effective logical queries that maximize retrieval efficiency, exploring strategies to reduce the number of LLM invocations, such as more efficient prompting, intermediate caching, or lightweight planning modules, could enhance the overall efficiency and practicality of our method in real-world applications.
Appendix A Dataset Details
GPQA Rein et al. (2024)    This dataset consists of physics, biology, and chemistry questions written by domain experts. We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance. Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve. Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions). Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
SuperGPQA (Team et al., 2025)    SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines. In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation. Each domain is further divided by three difficulty levels(easy, medium, and hard). To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42). Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
MMLU (Hendrycks et al., 2021; Wang et al., 2024b)    The MMLU benchmark covers a wide range of subjects across multiple domains. For out experiments, we focus on STEM and Professional categories. The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry. We specifically select two subsets: college chemistry from MMLU-STEM and professional chemistry from MMLU-Pro. These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
Appendix B Experiment Details
B.1 Baseline Details
PS Wang et al. (2023) Plan-and-Solve prompting is a direct reasoning strategy where the model first outlines a plan and then executes it step by step. We use the same prompt as the original paper.
Step-back (Direct reasoning) Zheng et al. (2024) Step-back prompting guides reasoning using abstracted principles derived from the original question. In our implementation, we follow the original Step-back paper’s prompting strategy to extract high-level principles necessary for solving the question. These extracted principles are then provided together with the original question as input to guide the model’s reasoning process.
CoT+RAG Lewis et al. (2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge. Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
Least-to-Most Zhou et al. (2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation. This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
Step-back (Decomposed reasoning) Zheng et al. (2024) We implement Step-back for decomposed reasoning by applying the Step-back prompting method to each subquestion in a decomposed reasoning framework. While the original Step-back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach. Specifically, after decomposing the original question into subquestions, we use the Step-back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
RAG (Decomposed reasoning) Liu et al. (2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query. The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
B.2 Retriever Configuration
Throughout all experiment, we retrieve top-10 documents per query.
B.3 RAISE
To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE.Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$. DPR similarity is computed as the inner product between L2-normalized query and passage embeddings. Higher scores indicate greater semantic similarity, with values closer to 1 representing stronger alignment between the query and passage.
We set $T=0.84$ for GPQA, SuperGPQA, and MMLU-Pro, which are composed of more challenging reasoning problems. For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
B.4 Algorithm
@@
B.5 Prompts
B.5.1 Baseline Prompts
Prompt for CoT You are solving a multiple choice question. Think step by step and show your reasoning clearly. At the end, state your answer in the format: "The final answer is (X)". Here, X must be the correct letter choice. Question: [Problem here] Answer Choices: [Answer choices here] Solution:
Prompt for Plan-and-Solve Let’s first understand the problem and devise a plan to solve the problem. Then, let’s carry out the plan and solve the problem step by step. Use the following format exactly: <Plan> Step 1: [state what you will do] Step 2: [state next reasoning step] … (as needed) <Solution> Step 1: [carry out Step 1] Step 2: [carry out Step 2] … (as needed) "The final answer is (X)." Here, X must be the correct letter choice. Question: [Problem here] Answer Choices: [Answer choices here] Solution:
Prompt for Step-back principle abstraction You are an expert at Science. You are given a Science problem. Your task is to extract the Science concepts and principles involved in solving the problem. What are the principles behind this question? End your response with "End of generation" after you answer the instructions. Question: [Subquestion here] Principles Involved:
Prompt for Step-back You are an expert at Science. You are given a Science problem and a set of principles involved in solving the problem. Solve the problem step by step by folowing the principles. At the end, state your answer in the format: "The final answer is (X)". Here, X must be the correct letter choice. Question: [Problem here] Principles: [Principles here] Answer Choices: [Answer choices here] Solution:
Prompt for hyDE query generation Generate a paragraph that answers the question. End your response with "End of generation" after you answer the instructions. Question: [Subquestion here] Explanation:
B.5.2 RAISE Prompts
Prompt for Problem Decomposition You are given a multiple-choice question. Break this problem into essential subquestions that directly help solve the original problem. Each subquestion MUST also include its search query. Each search query should reflect scientific or mathematical knowledge needed to answer the subquestion. STRICT FORMAT REQUIREMENTS: 1. For each subquestion, you MUST provide exactly two parts in this order: - The subquestion - A search query for that subquestion 2. Use EXACTLY this format for each subquestion: Subquestion 1: [your specific subquestion] Search Query for Subquestion 1: [Write a search query someone might realistically use to learn how to answer this subquestion] Question: [Problem here] Answer Choices: [Answer choices here]
Prompt for Logical Query Generation You are given a subquestion and a search query. The search query is a realistic phrase that someone might use to find knowledge or reasoning support to answer the subquestion. Your task is to anticipate what essential scientific or mathematical explanation the search result would contain, and write it concisely (2–3 sentences). Focus only on the core concept or principle that would help answer the subquestion. Avoid restating the subquestion, and do not include unrelated or overly general information. Subquestion: [Subquestion resulting from Problem Decomposition] Search Query: [Search query resulting from Problem Decomposition] Explanation:
Prompt for Solving Subquestions with Documents You are solving a multiple-choice question. Question is decomposed into several subquestions. You will be given: 1. The original multiple-choice question 2. Previous subquestions and their solutions (if any) 3. The current subquestion to solve 4. Documents that are relevant to the current subquestion Your task: - Carefully read the original question, any previous subquestion and their solutions, and the current subquestion - Use the information from the retrieved documents to solve the current subquestion - Also use your existing knowledge to solve the current subquestion - Your solution should be detailed and logically structured Documents: [Retrieved document] Question: [Problem here] Answer Choices: [Answer choices here] Previous subquestions and their solutions: [Previously generated subquestions and solutions] Current subquestion to solve: Subquestion [Step num]: [Subquestion] Subquestion [Step num] Solution:
Prompt for Generating Final Answer You are solving a multiple-choice question. Question is decomposed into several subquestions. Each subquestion has already been solved. Your task is to carefully read the original question and the several subquestion solutions, then use them to determine the final answer. Think step by step and then finish your answer with "The final answer is (X)" where X is the correct letter choice. Original Question: Question: [Problem here] Answer Choices: [Answer choices here] Subquestions and Solutions: [Generated stepwise subproblems and solutions] Final Solution:
Prompt for Evaluation with GPT You are given the following three items: - Original Problem: [Problem here] - Subquestion: [Subquestion here] - Retrieved Document: [Document here] Your task is to evaluate how helpful the retrieved document is for answering the subquestion. Please follow these instructions: - Do not just check if the topic is related. - Instead, check if the document includes information that helps someone reason through and solve the subquestion. - Focus on whether the document supports actual thinking or steps needed to get the answer. Give your final judgment using only one of the following ratings: - "No relevance at all" - does not have any domain similarity - "Superficially relevant" - has domain similarity (only superficially) but does not have any logical relevance to the subquestion. For example, the document might mention the same topic as the subquestion, but it does not provide any information that helps solve the subquestion. - "Partially relevant" - has domain similarity and has some logical relevance to the subquestion. For example, the document might provide some information that helps solve the subquestion, but it does not provide all the logical steps needed. - "Fully relevant" - has domain similarity and has almost all logical relevance to the subquestion. For example, the document provides enough relevant logical steps to solve the subquestion. Then explain your reasoning briefly. Output Format: Helpfulness Rating: <one of the 4 options above> Explanation: <your short explanation>
Appendix C Additional Results
@@
Table 5 provides the full subset results for each benchmark, complementing the averaged scores shown in Table 2. Across all datasets, RAISE consistently outperforms other retrieval-augmented baselines. The consistent improvement across subsets supports our central hypothesis that constructing queries enriched with logical intent is essential for retrieving logically relevant knoweldge.
Table 5: Comparison of decomposed reasoning strategies with retrieval. Subset-level results for each benchmark, complementing the averaged scores reported in the main text.
RAG Step-back +RAG HyDE RAISE
GPQA
Overall 45.95 43.43 48.48 51.01
SuperGPQA
science-hard 6.03 5.53 7.54 10.55
science-middle 15.08 15.58 13.07 19.60
engineering-hard 8.04 9.05 7.04 10.55
MMLU
(Pro) Chemistry 22.53 23.06 22.97 28.36
(STEM) College Physics 46.00 43.00 49.00 51.00

### v2

@@ 4.3 Analysis of RAISE
4.3.1 Impact of Problem Decomposition.
To further assess the importance of problem decomposition, we also evaluate a variant of our method that omits this step and directly performs logical query generation and retrieval without breaking the problem into subquestions, as shown in Figure 3. This version, referred to as RAISE-Direct, showed lower performance compared to the full version of RAISE. These results indicate that problem decomposition plays a critical role in guiding the retrieval process and structuring the reasoning pathway. This suggests that for complex reasoning problems, decomposing the question and retrieving logical knowledge tailored to each subquestion is more effective than retrieving once based on the original question alone. This is likely because different reasoning steps often require distinct pieces of information that may not be jointly retrievable from a single query.
4.3.2 Qualitative Evaluation of Retrieved Documents
We qualitatively demonstrate the examples when RAISE retrieves logically relevant documents compared to convential RAG in Table 3. While RAG often retrieves documents that are topically related yet fail to address the reasoning needs of the subquestion, RAISE consistently identifies documents that include essential scientific principles, mechanisms, or equations. For instance, in questions involving chemical reactions, RAISE surfaces materials that explain the specific reactivity or the retarded time calculation, whereas RAG may return vague definitions or unrelated economic concepts. These cases illustrate how RAISE’s retrieval is not only domain-aware but also aligned with the logical demands of solving complex scientific problems.
@@ 4.3.2 Qualitative Evaluation of Retrieved Documents
Additionally, since LLM-as-judge may struggle with assessing logical relevance in expert-level science problems like GPQA and SuperGPQA, we conducted human evaluations with 2–3 experts per question, including chemistry professors and PhDs.The results can be found in Appendix D.2. These human evaluation results further confirm that RAISE significantly outperforms other baselines in selecting logically relevant documents.
@@ 5 Related Works
LLMs for Scientific Reasoning.
Recent works have shown that LLMs can be applied for challenging scientific reasoning tasks. Unlike other domains, scientific reasoning requires not only step-by-step thinking, but also knowledge of specialized terminology and adaptation to continually evolving knowledge. Due to this challenging nature, many benchmarks have been proposed recently to tackle scientific reasoning with LLMs Rein et al. (2024); Zhong et al. (2025); Zhang et al. (2025b); Wang et al. (2024a); Jiang et al. (2024). Many works enhance scientific reasoning capabilities of LLMs through domain-specific training Taylor et al. (2022); Prabhakar et al. (2025); Zhang et al. (2024b), step-by-step reasoning Rueda et al. (2025); Rein et al. (2024), or retrieval of external knowledge or tools Ma et al. (2024); Zhong et al. (2025); Wellawatte et al. (2025); Li et al. (2025). Unlike previous works, we focus on applying step-by-step document retrieval from in-the-wild corpus without assuming access to well-curated and domain-specific corpus.
Step-wise Reasoning.
A growing body of research has shown that decomposing complex problems into structured intermediate steps can enhance the reasoning abilities of LLMs. An influential early approach, Chain-of-Thought prompting Wei et al. (2022), introduced explicit, sequential reasoning steps, making the model’s thought process more transparent and coherent. This inspired methods such as Plan-and-Solve Wang et al. (2023), which emphasizes high-level planning before answering, and Step-Back Prompting Zheng et al. (2024), which encourages abstraction by prompting the model to reflect before solving. Least-to-Most prompting Zhou et al. (2023) extends this by breaking down tasks into simpler subproblems, solved in increasing order of difficulty.
@@
Retrieval Augmented Generation.
Retrieval-Augmented Generation (RAG) Lewis et al. (2020b) was initially proposed to improve LLMs’ factual accuracy and knowledge by retrieving relevant external documents during generation Lewis et al. (2020b); Karpukhin et al. (2020); Izacard and Grave (2021); Khattab and Zaharia (2020).
@@ 6 Conclusion and Limitations
However, RAISE has some limitations and areas for future improvement. First, we used only DPR for retrieval. Incorporating recent models Gao et al. (2023b); Basu et al. (2024); Shi et al. (2023) may further improve performance. In addition, our study focuses on scientific reasoning. Applying RAISE to other domains such as legal, commonsense, or medical reasoning could help evaluate its generalizability. Moreover, RAISE requires multiple LLM calls during logical query generation. Reducing this cost through better prompting, caching, or lightweight planning could improve scalability and real-world usability.
Appendix A RAISE Algorithm
@@
Appendix B Dataset Details
GPQA
Rein et al. (2024) This dataset consists of physics, biology, and chemistry questions written by domain experts. We use GPQA diamond subset, which consist of 198 high-quality questions selected based on human performance. Specifically, this subset includes questions that both experts answer correctly while the majority of non-experts fail to solve. Each question typically demands multi-step reasoning, precise formula manipulation, and access to external scientific facts (e.g., physical constants, definitions). Due to its alignment with our target setting, GPQA serves as the primary evaluation benchmark throughout our experiments.
SuperGPQA
Team et al. (2025) SuperGPQA is a large-scale benchmark designed to evaluate graduate-level reasoning across 13 disciplines, 72 fields, and 285 graduate-level disciplines. In alignment with the scientific reasoning focus of our work, we select science and engineering domains for evaluation. Each domain is further divided by three difficulty levels(easy, medium, and hard). To reduce computational overhead while maintaining consistency, we randomly sample 199 questions per subset using a fixed seed (42). Specifically, our experiments include 199 examples each from science-hard, science-middle, and engineering-hard subsets.
MMLU
Hendrycks et al. (2021); Wang et al. (2024b) The MMLU benchmark covers a wide range of subjects across multiple domains. For out experiments, we focus on STEM and Professional categories. The STEM contains university-level science and engineering subjects such as college mathematics and computer science, while the Professional category covers specialized fields that typically require professional training or advanced education, including law, medicine, and chemistry. We specifically select three subsets: college chemistry from MMLU-STEM and professional chemistry and biology from MMLU-Pro. These subsets are chosen to evaluate our method’s ability to perform scientific reasoning in both academic and professional contexts involving complex domain knowledge.
Appendix C Experiment Details
C.1 Baseline Details
CoT+RAG
Lewis et al. (2020b) We implement CoT+RAG by combining Chain-of-Thought prompting with retrieval, where the model is prompted to solve the problem step-by-step while also leveraging external knowledge. Specifically, we provide the model with a CoT-style prompt encouraging step-by-step reasoning, alongside the original question and documents retrieved using the original question as the search query.
Least-to-Most
Zhou et al. (2023) Least-to-Most is a decomposed reasoning strategy that breaks down a complex problem into a sequence of simpler subquestions, which are then solved sequentially without retrieval augmentation. This subquestion decomposition pipeline serves as the foundational structure for other decomposed reasoning methods as well.
Step-Back
(Decomposed reasoning) Zheng et al. (2024) We implement Step-Back for decomposed reasoning by applying the Step-Back prompting method to each subquestion in a decomposed reasoning framework. While the original Step-Back paper does not cover the application of this method to decomposed subquestions, we extend it for a fair comparison with our approach. Specifically, after decomposing the original question into subquestions, we use the Step-Back prompting strategy to extract a high-level principle for each subquestion, and then provide the subquestion along with its corresponding principle to guide the model’s reasoning.
Least-to-Most+RAG
(Decomposed reasoning with RAG) Liu et al. (2025) We implement RAG by first decomposing the original problem into subquestions and then retrieving documents using each subquestion as a query. The retrieved documents are provided to the model along with the corresponding subquestion to support its reasoning.
Step-Back+RAG
(Decomposed reasoning with RAG) Zheng et al. (2024) We extend the Step-Back prompting strategy to a retrieval-augmented setting for fair comparison with our method. After decomposing the original question into subquestions, we generate a principle abstraction for each subquestion using Step-Back prompting, and use it as a query to retrieve evidence. The retrieved documents are then provided alongside the original subquestion to guide the model’s reasoning.
HyDE
(Decomposed reasoning with RAG) Gao et al. (2023a) We apply the HyDE approach to each subquestion in a decomposed reasoning framework. For each subquestion, the model first generates a hypothetical answer, which is then used as a query to retrieve supporting documents. The retrieved evidence, together with the subquestion, is provided to the model to support step-by-step reasoning.
C.2 Retriever Configuration
Throughout all experiment, we retrieve top-10 documents per query. To reduce the impact of potentially irrelevant documents by DPR, we apply a similarity threshold $T$ in RAISE. Specifically, we discard any retrieved passage whose DPR similarity score falls below $T$. DPR similarity is computed as the inner product between L2-normalized query and passage embeddings. Higher scores indicate greater semantic similarity, with values closer to 1 representing stronger alignment between the query and passage. We set $T=0.84$ for GPQA, SuperGPQA, and MMLU-Pro, which are composed of more challenging reasoning problems. For MMLU-STEM (college chemistry), we use a slightly lower threshold of $T=0.80$, considering that the questions are generally simpler than those in other datasets.
Appendix D Additional Results
D.1 Further Analysis on GPQA
@@
D.2 Human Evaluation of Retrieved Documents
Figure 6: Human Evaluation of the Logical Relevance of Retrieved Documents Aligned with the results from the LLM-as-a-judge evaluation of logical relevancy, RAISE shows a higher proportion of logically relevant documents and a lower proportion of irrelevant or superficially relevant ones.
Previously, in Section 4.3.2, we evaluated the logical relevancy of documents retrieved by the baseline methods and RAISE using the LLM-as-a-judge approach. However, since our domain includes complex, expert-level questions, and LLM-based evaluations may overlook domain-specific reasoning and often rely on surface-level features, we supplemented our analysis with a small-scale human evaluation of 20 subquestion–document pairs. Each pair was assessed by at least three annotators, including Ph.D. students and a faculty member in chemistry, with the method provenance concealed to maintain objectivity. The results are presented in Figure 6.
As also discussed in the LLM-as-a-judge results, the human evaluation indicates that RAISE produces significantly fewer irrelevant documents compared to all other methods, while achieving the highest proportion of logically relevant documents. Although limited in scale due to time and cost constraints, we believe this evaluation provides meaningful human validation of RAISE’s effectiveness and serves as a valuable complement to the LLM-based assessments.
Appendix E Prompts
E.1 Baseline Prompts
You are solving a multiple choice question. Think step by step and show your reasoning clearly.
At the end, state your answer in the format: "The final answer is (X)".
Here, X must be the correct letter choice.
Question: [Problem here]
Answer Choices: [Answer choices here]
Solution:
Figure 7: Prompt for CoT
You are an expert at Science. You are given a Science problem.
Your task is to extract the Science concepts and principles involved in solving the problem.
What are the principles behind this question?
End your response with "End of generation" after you answer the instructions.
Question: [Subquestion here]
Principles Involved:
Figure 8: Prompt for Step-Back Principle Abstraction
You are an expert at Science. You are given a Science problem and a set of principles involved in solving the problem.
Solve the problem step by step by following the principles.
At the end, state your answer in the format: "The final answer is (X)".
Here, X must be the correct letter choice.
Question: [Problem here]
Principles: [Principles here]
Answer Choices: [Answer choices here]
Solution:
Figure 9: Prompt for Step-Back
Generate a paragraph that answers the question.
End your response with "End of generation" after you answer the instructions.
Question: [Subquestion here]
Explanation:
Figure 10: Prompt for HyDE Query Generation
E.2 RAISE Prompts
You are given a multiple-choice question.
Break this problem into essential subquestions that directly help solve the original problem.
Each subquestion MUST also include its search query.
Each search query should reflect scientific or mathematical knowledge needed to answer the subquestion.
STRICT FORMAT REQUIREMENTS:
1. For each subquestion, you MUST provide exactly two parts in this order:
- The subquestion
- A search query for that subquestion
2. Use EXACTLY this format for each subquestion:
Subquestion 1: [your specific subquestion]
Search Query for Subquestion 1: [Write a search query someone might realistically use to learn how to answer this subquestion]
Question: [Problem here]
Answer Choices: [Answer choices here]
Figure 11: Prompt for Problem Decomposition
You are given a subquestion and a search query.
The search query is a realistic phrase that someone might use to find knowledge or reasoning support to answer the subquestion.
Your task is to anticipate what essential scientific or mathematical explanation the search result would contain, and write it concisely (2–3 sentences).
Focus only on the core concept or principle that would help answer the subquestion.
Avoid restating the subquestion, and do not include unrelated or overly general information.
Subquestion: [Subquestion resulting from Problem Decomposition]
Search Query: [Search query resulting from Problem Decomposition]
Explanation:
Figure 12: Prompt for Logical Query Generation
You are solving a multiple-choice question. The question is decomposed into several subquestions. You will be given:
 1. The original multiple-choice question
 2. Previous subquestions and their solutions (if any)
 3. The current subquestion to solve
 4. Documents that are relevant to the current subquestion
Your task:
- Carefully read the original question, any previous subquestions and their solutions, and the current subquestion.
- Use the information from the retrieved documents to solve the current subquestion.
- Also use your existing knowledge to solve the current subquestion.
- Your solution should be detailed and logically structured.
Documents: [Retrieved document]
Question: [Problem here]
Answer Choices: [Answer choices here]
Previous subquestions and their solutions:
[Previously generated subquestions and solutions]
Current subquestion to solve:
Subquestion [Step num]: [Subquestion]
Subquestion [Step num] Solution:
Figure 13: Prompt for Solving Subquestions with Documents
You are solving a multiple-choice question. The question is decomposed into several subquestions. Each subquestion has already been solved. Your task is to carefully read the original question and the several subquestion solutions, then use them to determine the final answer. Think step by step and then finish your answer with "The final answer is (X)" where X is the correct letter choice.
Original Question:
Question: [Problem here]
Answer Choices: [Answer choices here]
Subquestions and Solutions:
[Generated stepwise subproblems and solutions]
Final Solution:
Figure 14: Prompt for Generating Final Answer
You are given the following three items:
- Original Problem: [Problem here]
- Subquestion: [Subquestion here]
- Retrieved Document: [Document here]
Your task is to evaluate how helpful the retrieved document is for answering the subquestion.
Please follow these instructions:
- Do not just check if the topic is related.
- Instead, check if the document includes information that helps someone reason through and solve the subquestion.
- Focus on whether the document supports actual thinking or steps needed to get the answer.
Give your final judgment using only one of the following ratings:
- "No relevance at all" – does not have any domain similarity
- "Superficially relevant" – has domain similarity (only superficially) but does not have any logical relevance to the subquestion. For example, the document might mention the same topic as the subquestion, but it does not provide any information that helps solve the subquestion.
- "Partially relevant" – has domain similarity and has some logical relevance to the subquestion. For example, the document might provide some information that helps solve the subquestion, but it does not provide all the logical steps needed.
- "Fully relevant" – has domain similarity and has almost all logical relevance to the subquestion. For example, the document provides enough relevant logical steps to solve the subquestion.
Then explain your reasoning briefly.
Output Format:
Helpfulness Rating: <one of the 4 options above>
Explanation: <your short explanation>
Figure 15: Prompt for Evaluation with GPT